A 64 -centimeter-long board is cut into 3 pieces. The longest piece is twice as long as the shortest piece. The other piece piece is 4 cm longer that the shortest piece. How long is each piece?

Answer by ReadingBoosters(3246)

x = longest piece
y = shortest piece
z = 3rd piece
x + y + z = 64cm
x = 2y
z = y + 4
Substitute the second and third equations into the first equation
2y + y + y + 4 = 64
Combine like terms
4y + 4 = 64
4y + 4 - 4 = 64 - 4
4y = 60
y = 15 cm is the shortest piece
x = 2y = 2(15) = 30cm is the longest piece
z = y + 4 = 15 + 4 = 19cm is the third piece